# Formalizing the Safety, Security, and Functional Properties of Agentic AI Systems

### v1

@@ IV-A Constructing the Host Agent Model
Definition of the Host Agent Model ($\mathcal{H}$). We define the host agent model $\mathcal{H}$ as a tuple: $\mathcal{H}=\left(\mathcal{A},\mathcal{E},\mathcal{T},\mathcal{R},\mathcal{C},\mathcal{O},\mathcal{C_{L}},\mathcal{S}_{\mathcal{H}}\right)$ where the components are defined as follows:
@@
  • $\mathcal{E}$: The set of all External Entities (EEs), such that $\mathcal{A}\subseteq\mathcal{E}$, which includes functional tools (MCP Servers).
@@
  • $\mathcal{C}$: The Host Agent Core (HAC), responsible for intent resolution $I_{U}$, $\mathcal{C}:\mathcal{T}\times\mathcal{S}_{SM}\to\mathcal{I}_{U}$.
  • $\mathcal{O}$: The Orchestrator, responsible for task decomposition, execution management, and final aggregation.
  • $\mathcal{C_{L}}$: The Communication Layer (CL), responsible for secure, reliable and protocol-agnostic communication with External Entities in $\mathcal{E}$.
@@
The Orchestrator’s core function is task decomposition, mapping a resolved intent $I_{U}$ and available entities $\mathcal{E}$ to a Directed Acyclic Graph (DAG) of sub-tasks $\mathcal{D}$: $\mathcal{O}_{\mathrm{decomp}}:\mathcal{I}_{U}\times\mathcal{P}(\mathcal{E})\to\mathcal{D}$. The execution management function, $\mathcal{O}_{\mathrm{exec}}$, maps the DAG through sub-task invocations (via $\mathcal{D}$) to the final results: $\mathcal{O}_{\mathrm{exec}}:\mathcal{D}\to\mathcal{F}_{\mathrm{results}}$
@@
The host agent model provides the foundational state space $\mathcal{S}_{\mathcal{H}}$ and critical functions ($\mathcal{C},\mathcal{O}$) to define the liveness, safety, completeness, fairness, and reachability properties in Table I.
IV-A1 Functional Architecture and Execution Sequence
We detail the operation of HA by referencing specific components and their interactions ( A- E), and execution steps ( 1- 8). This defines the role of each component and the data flow from the initial user prompt to the final response.
  • Session Manager (SM): The SM maintains dialogue context across multi-turn interactions and manages user-stored credentials for secure access to authenticated services. We integrate the SM based on evidence that contextual management improves decision-making and reasoning in LLM-based systems [7, 37, 38].
  • Dialogue Manager (DM): The DM handles conversation flow. It works with the LLM to clarify user intent, particularly when a request is ambiguous, eliciting necessary information before execution proceeds ( 1’).
The HA’s context-aware reasoning resolves the user intent ($\mathcal{I}_{U}$), captured formally as:
$\text{HAC}(Req_{U},\text{State}_{SM})\to\mathcal{I}_{U}$
where $\text{State}_{SM}$ represents the current dialogue and credential state maintained by the SM. Following intent resolution, the HAC initiates the subsequent task phases.
IV-B Constructing the Task Lifecycle Model
Figure 2: The task lifecycle model that details sub-task state transitions from creation to termination. The lifecycle begins in the CREATED state, proceeds through AWAITING DEPENDENCY (if prerequisites exist), and moves to READY. Execution transitions either directly to IN PROGRESS (internal handling) or via DISPATCHING (external delegation). Successful execution leads to COMPLETED. Failures transition to FAILED, enabling recovery via RETRY SCHEDULED or FALLBACK SELECTED, before terminating in ERROR. Sub-tasks may be CANCELED at any stage.
Registry - B. The Registry maintains a dynamic collection of external entities (EEs)–including autonomous AI agents (i.e., A2A servers [5]), functional tools (i.e., MCP servers [4]), and other service endpoints (contained in E). These entities are registered to expose capabilities leveraged during task execution. Each EE requires a capability profile ($\mathcal{P}(\mathcal{EE}_{info}\times\mathcal{M}_{API})$), which details its skills ($\mathcal{EE}_{info}$) and associated API metadata ($\mathcal{M}_{API}$) necessary for invocation. The Registry supports entity registration, efficient querying, and deregistration to enable system extensibility.
Following intent resolution, the HAC queries the Registry ( 2) to identify available EEs suitable for fulfilling the request. We formalize EEs discovery and registration as:
$\mathsf{\mathcal{R}.discover}(EE)\to\mathcal{P}(\mathcal{EE}_{info}\times\mathcal{M}_{API})$
$\mathsf{\mathcal{R}.register}(\mathcal{P}(\mathcal{EE}_{info}\times\mathcal{M}_{API}))\to\text{Success}\mid\text{Failure}$
Orchestrator - C Upon receiving the resolved user intent ($\mathcal{I}_{U}$) and a set of relevant EEs from the Registry, the LLM constructs a structured execution plan. This plan decomposes the intent into interrelated sub-tasks organized as a Directed Acyclic Graph (DAG) ( 3), which is defined as:
$\mathcal{D}=(\mathcal{V},\mathcal{S})$
where $v\in\mathcal{V}$ is a sub-task node, and $\mathcal{S}$ is the set of directed edges indicating precedence constraints or data dependencies.
The DAG structure supports parallelism, scheduling, and deterministic execution by capturing dependencies. A directed edge $(v_{i},v_{j})\in\mathcal{S}$ implies that sub-task $v_{j}$ executes only after sub-task $v_{i}$ successfully completes.
The Orchestrator manages the DAG’s execution. It schedules and invokes sub-tasks ( 4) according to the dependency structure to ensure prerequisites are satisfied. After external entities execute, the Orchestrator collects results ( 6) and aggregates them ( 7) by applying predefined aggregation functions, merging partial outputs into a unified data structure that satisfies the original task requirements. This execution process is primarily defined by:
$D\>:=\>\mathsf{LLM.Build\_Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\})$
$\mathsf{O.execute}(D)\to\text{results}$
$\bigwedge_{i=1}^{n=|\mathcal{D}|}\mathsf{O.aggregate}(sub\_task_{1},\dots,sub\_task_{n})$
Communication Layer (CL) - D The CL ensures secure, reliable, and protocol-agnostic communication between the HA and EEs in E. It supports multiple communication protocols, such as REST and gRPC, where a protocol defines the communication mechanism and serialization format (e.g., gRPC with Protocol Buffers).
The CL routes sub-tasks for external invocation ( 5) and handles communication using the appropriate protocol and payload. The primary interface provided is:
$\mathsf{CL.invoke}(\text{EE},\text{protocol},\text{payload})\to\text{response}$
The payload contains the sub-task description or API call parameters. The response includes results, error codes, or computation artifacts, which the Orchestrator receives. Finally, the HAC generates the final response ($Resp_{H}$) through the LLM and delivers it to the user via the DM ( 8).
This overall sequence demonstrates the system’s intrinsic modularity, its reliance on secure communication channels handled by the CL, and the advanced task orchestration capability inherent in the Orchestrator.
Building upon the HA model, we introduce the task lifecycle model to provide a fine-grained perspective on how the HA manages and executes sub-tasks. While the HA model (Figure 1) outlines the system architecture and orchestration, the task lifecycle model (Figure 2) adds a layer of abstraction by detailing the states and transitions a sub-task undergoes from its origin to completion or failure.
Definition of the Task Lifecycle Model ($\mathcal{L}$). The task lifecycle model $\mathcal{L}$ defines the state space and transition function for a single sub-task $t$, where $t$ is a constituent node in the host agent’s decomposition graph $D$. $\mathcal{L}=\left(\mathcal{S}_{t},s_{0},\mathcal{E}_{t},\delta\right)$ where:
    $\mathcal{S}_{t}={\footnotesize{\left\{\begin{array}[]{@{}l@{}}\texttt{CREATED},\texttt{AWAITING DEPENDENCY},\texttt{READY},\\ \texttt{DISPATCHING},\texttt{IN PROGRESS},\texttt{COMPLETED},\\ \texttt{FAILED},\texttt{RETRY SCHEDULED},\\ \texttt{FALLBACK SELECTED},\texttt{CANCELED},\texttt{ERROR}\end{array}\right\}}}$
  • $s_{0}$: The initial state, $s_{0}=\texttt{CREATED}$.
This model’s state space $\mathcal{S}_{t}$ and transition function $\delta$ are fundamental to verifying task execution integrity. They provide the basis for the safety invariants (e.g., sequential transition checks) and liveness and fairness properties (e.g., eventual termination/progression to READY state), detailed in Section V.
Task and Sub-task Dynamics. In a lifecycle model, a task represents the high-level user request that serves as the core unit of work fulfilled by the HA. A sub-task is a lower-level, constituent component into which the HAC decomposes the main task. Each sub-task has its own lifecycle, tracked within the parent task’s context. This lifecycle with a defined sequence of states supports fine-grained control for asynchronous execution across distributed EEs.
TABLE I: Specification of Host Agent Formal Model Properties ($\mathtt{HP_{1}}$$\mathtt{HP_{17}}$). These temporal logic invariants (CTL/LTL) define the requirements for the overall system orchestration. The properties verify the integrity of the delegation workflow, adherence to security policies (e.g., Trust Anchor safety), and reliable progression toward task completion across all core components (Host Agent Core, Registry, Orchestrator, and Communication Layer).
@@ IV-B Constructing the Task Lifecycle Model
$AG(\mathcal{I}_{U}\rightarrow AF\>\mathsf{LLM.Build\_Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\}))$
@@
$AG(\mathsf{LLM.Build\_Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\}))\rightarrow\left(\bigwedge_{i=1}^{n=|D|}\mathsf{AF}\>(\mathsf{CL.invoke}(EE,protocol,sub\_task_{i}))\right)$
@@
$\mathsf{AG}\>(\mathsf{CL.invoke}(EE,protocol,sub\_task))\rightarrow AF\>(\mathsf{CL.return\_result}(sub\_task))$
@@
$\bigwedge_{i=1}^{n=|D|}\mathsf{AG}(\mathsf{CL.return\_result}(sub\_task_{i}))\rightarrow\mathsf{AF}(\mathsf{O.aggregate}(sub\_task_{1},\dots,sub\_task_{n}))$
@@
$AG(\mathcal{R}.entities\rightarrow AX(\mathsf{LLM.Build\_Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\})))$
@@
$AG(\mathsf{CL.invoke}(EE,protocol,sub\_task)\rightarrow sub\_task\in D)$
$\mathtt{HP_{9}}$ Task invocation is strictly conditional on the EE having completed the system’s predefined validation process and possessing a minimum status of reliability.
$AG\>(\mathsf{CL.invoke}(\text{EE},\text{protocol},\text{payload})\rightarrow VM(\text{EE}))$
@@
$\bigwedge_{i=1}^{n=|D|}AG(\mathsf{CL.invoke}(EE,protocol,sub\_task_{i}))\rightarrow dependencies[sub\_task_{i}]=\varnothing$
@@
$Resp_{H}\rightarrow\left(\bigwedge_{i=1}^{n=|D|}\mathsf{AG}(\mathsf{CL.invoke}(EE,protocol,sub\_task_{i}))\right)$
@@
$AG(Req_{U}\rightarrow EX(\mathsf{LLM.Build\_Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\}))\lor\text{Clarify\_Intent}))$
@@
$\mathrm{FAIRNESS}(\mathsf{Agent\_RPC\_Calls})$
@@
$\mathrm{FAIRNESS}(\mathsf{JSON\_RPC\_Calls})$
$\mathtt{HP_{15}}$ Every sub-task invocation eventually yields a result.
$\mathrm{FAIRNESS}(\mathsf{CL.return\_result}(sub\_task))$
@@
$\mathtt{HP_{16}}$ It is always possible to reach a state in which the Host Agent replies to the user.
@@
$\mathtt{HP_{17}}$ It is always possible to reach a state in which the LLM builds a Task DAG.
$EF(\mathsf{LLM.Build\_Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\}))$
TABLE II: Specification of Task Lifecycle Formal Model Properties ($\mathtt{TL_{1}}$$\mathtt{TL_{14}}$). These temporal logic invariants (CTL/LTL) define the requirements for robust sub-task execution. The properties verify the integrity of state transitions, dependency satisfaction, and reliable failure handling mechanisms.
@@
$\mathtt{TL_{1}}$ Every sub-task that is CREATED eventually terminates in one of the states: COMPLETED, ERROR, or CANCELED.
$AG(state=\texttt{CREATED}\rightarrow AF(state=\texttt{COMPLETED}\lor state=\texttt{ERROR}\lor state=\texttt{CANCELED}))$
$\mathtt{TL_{2}}$ A sub-task in READY that requires an external entity will eventually transition to DISPATCHING.
@@
$\mathtt{TL_{3}}$ A sub-task in FALLBACK SELECTED eventually transitions to either DISPATCHING, CANCELED, or FAILED.
$AG(state=\texttt{FALLBACK SELECTED}\rightarrow AF(state=\texttt{DISPATCHING}\lor state=\texttt{CANCELED}\lor state=\texttt{FAILED}))$
$\mathtt{TL_{4}}$ A sub-task in DISPATCHING will eventually reach the IN PROGRESS state.
@@
$\mathtt{TL_{5}}$ A sub-task may only enter DISPATCHING if its previous state was READY.
$G((state=\texttt{DISPATCHING})\rightarrow(previous\_state=\texttt{READY}))$
$\mathtt{TL_{6}}$ A sub-task may only enter COMPLETED if it was previously IN PROGRESS.
$G((state=\texttt{COMPLETED})\rightarrow(previous\_state=\texttt{IN PROGRESS}))$
$\mathtt{TL_{7}}$ Once a sub-task enters ERROR, it remains there and cannot return to an active processing state.
$AG((state=\texttt{ERROR})\rightarrow AG(state=\texttt{ERROR}))$
$\mathtt{TL_{8}}$ A sub-task may only enter RETRY SCHEDULED if its previous state was FAILED.
@@
$\mathtt{TL_{9}}$ A sub-task that enters CANCELED remains permanently in that state.
$AG((state=\texttt{CANCELED})\rightarrow AG(state=\texttt{CANCELED}))$
$\mathtt{TL_{10}}$ A sub-task cannot remain indefinitely in AWAITING DEPENDENCY; it must eventually progress to another state.
Fairness
$\mathtt{TL_{11}}$ A sub-task in AWAITING DEPENDENCY with all dependencies satisfied eventually transitions to READY.
Specific Flow/Path Properties
$\mathtt{TL_{12}}$ A sub-task in FAILED with no available fallbacks transitions either to RETRY SCHEDULED or directly to ERROR.
$AG((state=\texttt{FAILED}\land\neg has\_fallbacks)\rightarrow(AX(state=\texttt{RETRY SCHEDULED}\lor AX(state=\texttt{ERROR})))$
$\mathtt{TL_{13}}$ A sub-task in FAILED with available fallbacks transitions to one of: RETRY SCHEDULED, FALLBACK SELECTED, or ERROR.
$AG((state=\texttt{FAILED}\land has\_fallbacks)\rightarrow(AX(state=\texttt{RETRY SCHEDULED}\lor AX(state=\texttt{FALLBACK SELECTED})\lor AX(state=\texttt{ERROR}))))$
@@
V Property Specification
The formalization of the HA and Task Lifecycle models provides the foundation for analyzing and verifying critical system properties. In this context, formal properties are verifiable statements about the system’s execution paths, used to ensure the safety, security, and functionality of an Agentic AI system’s behavior. We structure these properties into two distinct sets: those captured by the HA model (overall orchestration, Table I), and those governing the task lifecycle model (sub-task state transitions, Table II).
These properties do not operate in isolation; they form a tightly coupled system of interdependent guarantees necessary for correct and reliable function in a multi-AI agent environment. Both sets of properties are expressed using temporal logics, Computation Tree Logic (CTL) and Linear Temporal Logic (LTL). These logics allow for the precise, unambiguous definition of desired system behavior over time and execution paths.
V-A Property Design Principles and Definitions
Both models include a set of property categories, each containing multiple specific properties, to establish verifiable guarantees. The core categories used in both tables are liveness, safety, ordering, completeness, fairness, and reachability. These properties ensure the overall system makes progress, avoids undesirable states, achieves its goals, and handles reliable resource allocation.
Liveness. Liveness ensures that an Agentic AI system will eventually make progress toward a goal. This property guarantees that, despite asynchronous actions, inter-agent dependencies, or potential conflicts, the system avoids global deadlock or infinite starvation. This is critical in these systems, where a lack of progress in one agent can propagate and result in a cascading stall of the entire system.
  • Example (Host Agent Model, $\mathtt{HP_{1}}$): Every user prompt ($Req_{U}$) must eventually receive a final response ($Resp_{H}$) from the host agent. This is expressed in temporal logic as $AG(Req_{U}\to AF\ Resp_{H})$.
  • Example (Task Lifecycle Model, $\mathtt{TL_{1}}$): A sub-task that is CREATED must eventually terminate in a defined end state (COMPLETED, ERROR, or CANCELED).

### v2

@@ 4.1 Constructing the Host Agent Model
Definition of the Host Agent Model ($\mathcal{H}$). We define the HA model $\mathcal{H}$ as a tuple: $\mathcal{H}=\left(\mathcal{A},\mathcal{E},\mathcal{T},\mathcal{R},\mathcal{C},\mathcal{O},\mathcal{C_{L}},\mathcal{S}_{\mathcal{H}}\right)$ where the components are defined as follows:
@@
  • $\mathcal{E}$: The set of all External Entities (EEs), such that $\mathcal{A}\subseteq\mathcal{E}$, which includes functional tools (MCP).
@@
  • $\mathcal{C}$: Host Agent Core (HAC), responsible for intent resolution $I_{U}$, $\mathcal{C}:Req_{U}\times\mathcal{S}_{SM}\to\mathcal{I}_{U}$.
  • $\mathcal{O}$: Orchestrator, responsible for task decomposition, execution management, and aggregation.
  • $\mathcal{C_{L}}$: Communication Layer (CL), responsible for secure, reliable, and protocol-agnostic communication with External Entities in $\mathcal{E}$.
@@
The Orchestrator’s core function is task decomposition, mapping a resolved intent $I_{U}$ and available entities $\mathcal{E}$ to a Directed Acyclic Graph (DAG) of sub-tasks $\mathcal{D}$, $\mathcal{O}_{\mathrm{decomp}}:\mathcal{I}_{U}\times\mathcal{R}(\mathcal{E})\to\mathcal{D}$. The execution management function, $\mathcal{O}_{\mathrm{exec}}$, maps the DAG through sub-task invocations (via $\mathcal{D}$) to the final results to be aggregated: $\mathcal{O}_{\mathrm{exec}}:\mathcal{D}\to\mathcal{F}_{\mathrm{results}}$.
@@
The HA model provides the state space $\mathcal{S}_{\mathcal{H}}$ and critical functions ($\mathcal{C},\mathcal{O}$) to define the liveness, safety, completeness, fairness, and reachability properties in Table 1.
Table 1: Specification of Host Agent formal model properties ($\mathtt{HP_{1}}$$\mathtt{HP_{16}}$).
@@
$AG(\mathcal{I}_{U}\rightarrow AF\>\mathsf{LLM.Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\}))$
@@
$AG(\mathsf{LLM.Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\})\rightarrow\forall i\in D:\mathsf{AF}\>(\mathsf{CL.invoke}(EE,prot,sub\_task_{i})))$
@@
$\mathsf{AG}\>(\mathsf{CL.invoke}(EE,prot,sub\_task)\rightarrow AF\>(\mathsf{CL.return\_result}(sub\_task)))$
@@
$AG((\forall i\in D:\mathsf{HasResult}(sub\_task_{i}))\rightarrow\mathsf{AF}(\mathsf{O.aggregate}(sub\_task_{1},\dots,sub\_task_{n})))$
@@
$AG\Big(\mathsf{LLM.Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\})\rightarrow\forall e\in\{\mathcal{EE}_{info}\}:\mathsf{is\_valid}(\mathsf{\mathcal{R}.discover}(e))\Big)$
@@
$AG(\mathsf{CL.invoke}(EE,prot,sub\_task)\rightarrow sub\_task\in D)$
$\mathtt{HP_{9}}$ Task invocation is strictly conditional on the EE having completed the system’s predefined validation process.
$AG\>(\mathsf{CL.invoke}(\text{EE},\textit{prot},\textit{payload})\rightarrow VM(\text{EE}))$
@@
$AG\Big(\forall i\in D:\mathsf{CL.invoke}(EE,prot,sub\_task_{i})\rightarrow\forall p\in\mathsf{parents}(sub\_task_{i}):\mathsf{Completed}(p)\Big)$
@@
$AG\Big(Resp_{H}\rightarrow(\forall i\in D:\mathsf{WasInvoked}(sub\_task_{i}))\Big)$
@@
$AG(Req_{U}\rightarrow AF(\mathsf{LLM.Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\})\lor\text{Clarify\_Intent}))$
@@
$AG(\mathsf{A2A\_Call}\rightarrow AF(\mathsf{A2A\_Response}))$
@@
$AG(\mathsf{MCP\_Call}\rightarrow AF(\mathsf{MCP\_Response}))$
@@
$\mathtt{HP_{15}}$ It is always possible to reach a state in which the Host Agent replies to the user.
@@
$\mathtt{HP_{16}}$ It is always possible to reach a state in which the LLM builds a Task DAG.
$EF(\mathsf{LLM.Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\}))$
4.1.1 Functional Architecture and Execution Sequence.
We detail the operation of HA by referencing specific components and their interactions ( A- E), as well as execution steps ( 1- 8), as depicted in Figure 1. This defines the role of each component and the data flow from the initial user prompt to the final response.
  • Session Manager (SM): The SM maintains dialogue context across multi-turn interactions and manages user-stored credentials for secure access to authenticated services. We integrate the SM based on evidence that contextual management improves decision-making and reasoning in LLM-based systems (Han et al., 2025; Yao et al., 2023; Li and Qiu, 2023).
  • Dialogue Manager (DM): The DM handles conversation flow, working with the LLM to clarify ambiguous user intent and elicit necessary information before execution proceeds ( 1’).
The HA’s context-aware reasoning resolves the user intent ($\mathcal{I}_{U}$), captured formally as $\text{HAC}(Req_{U},\text{State}_{SM})\to\mathcal{I}_{U}$, where $\text{State}_{SM}$ represents the current dialogue and credential state maintained by the SM. Following intent resolution, the HAC initiates the subsequent task phases.
Registry - B. The Registry maintains a dynamic collection of EEs–including autonomous AI agents (i.e., A2A server), functional tools (i.e., MCP servers), and other service endpoints (contained in E). These entities are registered to expose capabilities leveraged during task execution. Each EE requires a capability profile ($\mathcal{P}(\mathcal{EE}_{info}\times\mathcal{M}_{API})$), which details its skills ($\mathcal{EE}_{info}$) and associated API metadata ($\mathcal{M}_{API}$) necessary for invocation. The Registry supports entity registration, efficient querying, and deregistration to enable system extensibility. Post-resolution, the HAC leverages the Registry ( 2) to discover suitable EEs. Discovery is defined as $\mathsf{\mathcal{R}.discover}(EE)\to\mathcal{P}(\mathcal{EE}_{info}\times\mathcal{M}_{API})$, and entity registration as $\mathsf{\mathcal{R}.register}(\mathcal{P}(\mathcal{EE}_{info}\times\mathcal{M}_{API}))\to\textit{Success}\mid\textit{Failure}$.
Orchestrator - C. Using intent $\mathcal{I}_{U}$ and Registry entities, the LLM builds a DAG $\mathcal{D}=(\mathcal{V},\mathcal{S})$ ( 3) where nodes $\mathcal{V}$ are interrelated sub-tasks and edges $\mathcal{S}$ enforce precedence constraints or data dependencies. An edge $(v_{i},v_{j})$ indicates $v_{j}$ must await $v_{i}$’s completion. The Orchestrator schedules ( 4), collects ( 6), and aggregates ( 7) results. This process is defined by $D:=\mathsf{LLM.Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\})$ and subsequent execution $\mathsf{O.execute}(D)$ where final output is aggregated $\forall i\in D:\mathsf{O.aggregate}(v_{1},\dots,v_{n})$.
Communication Layer (CL) - D. The CL ensures secure, protocol-agnostic interaction with EEs ( E), supporting standards such as REST and gRPC. It routes sub-tasks for invocation ( 5) via $\mathsf{CL.invoke}(\text{EE},\textit{prot},\textit{payload})\to\textit{resp}$. The payload encapsulates API parameters, while resp returns results or errors to the Orchestrator. Finally, the HAC synthesizes the user output ($Resp_{H}$) via the LLM, delivering it through the DM ( 8).
Figure 2: Sub-task state transitions including execution, delegation, recovery, and termination.
4.2 Constructing the Task Lifecycle Model
Building upon the HA model, we introduce the task lifecycle model (Figure 2). It abstracts the sub-task lifecycle by detailing state transitions from origin to termination.
Definition of the Task Lifecycle Model ($\mathcal{L}$). The task lifecycle model $\mathcal{L}$ defines the state space and transition function for a single sub-task $t$, where $t$ is a constituent node in the host agent’s decomposition graph $D$. Task lifecycle model $\mathcal{L}=\left(\mathcal{S}_{t},s_{0},\mathcal{E}_{t},\delta\right)$:
[…]
  • $s_{0}$: The initial state, $s_{0}=\texttt{CREATED}$.
The state space $\mathcal{S}_{t}$ and transition function $\delta$ are fundamental to verifying task execution integrity. They provide the basis for the safety invariants (e.g., sequential transition checks) and liveness and fairness properties (e.g., eventual termination/progression to READY state), detailed in Section 5.
Task and Sub-task Dynamics. In the lifecycle model, a task represents the high-level user request that serves as the core unit of work fulfilled by the HA. A sub-task is a lower-level, constituent component into which the HAC decomposes the main task. Each sub-task has its own lifecycle, tracked within the parent task’s context. This lifecycle with a defined sequence of states supports fine-grained control for asynchronous execution across distributed EEs.
Table 2: Specification of Task Lifecycle formal model properties.
@@ 4.2 Constructing the Task Lifecycle Model
$\mathtt{TL_{1}}$ Every CREATED sub-task eventually terminates in COMPLETED, ERROR, or CANCELED.
$AG(state=\texttt{CREATED}\rightarrow AF(state\in\{\texttt{COMPLETED},\texttt{ERROR},\texttt{CANCELED}\}))$
$\mathtt{TL_{2}}$ A sub-task in READY that requires an external entity eventually transitions to DISPATCHING.
@@
$\mathtt{TL_{3}}$ A sub-task in FALLBACK SELECTED eventually transitions to DISPATCHING, CANCELED, or FAILED.
$AG(state=\texttt{FALLBACK SELECTED}\rightarrow AF(state\in\{\texttt{DISPATCHING},\texttt{CANCELED},\texttt{FAILED}\}))$
$\mathtt{TL_{4}}$ A sub-task in DISPATCHING eventually reaches the IN PROGRESS state.
@@
$\mathtt{TL_{5}}$ A sub-task cannot remain indefinitely in AWAITING DEPENDENCY.
$\mathtt{TL_{6}}$ A sub-task in AWAITING DEPENDENCY with satisfied dependencies eventually transitions to READY.
@@
$\mathtt{TL_{7}}$ A sub-task enters DISPATCHING only from READY, FALLBACK SELECTED, or RETRY SCHEDULED.
$AG((state=\texttt{DISPATCHING})\rightarrow previous\_state\in\{\texttt{READY},\texttt{FALLBACK SELECTED},\texttt{RETRY SCHEDULED}\})$
$\mathtt{TL_{8}}$ A sub-task may only enter COMPLETED if it was previously IN PROGRESS.
$AG((state=\texttt{COMPLETED})\rightarrow(previous\_state=\texttt{IN PROGRESS}))$
$\mathtt{TL_{9}}$ Once a sub-task enters ERROR, it remains there permanently.
$AG((state=\texttt{ERROR})\rightarrow AX(state=\texttt{ERROR}))$
$\mathtt{TL_{10}}$ A sub-task may only enter RETRY SCHEDULED if its previous state was FAILED.
@@
$\mathtt{TL_{11}}$ A sub-task that enters CANCELED remains permanently in that state.
$AG((state=\texttt{CANCELED})\rightarrow AX(state=\texttt{CANCELED}))$
Transition Constraints
$\mathtt{TL_{12}}$ A FAILED sub-task with no fallbacks transitions to RETRY SCHEDULED, ERROR or is CANCELED.
$AG((state=\texttt{FAILED}\land\neg has\_fallbacks)\rightarrow AX(state\in\{\texttt{RETRY SCHEDULED},\texttt{ERROR},\texttt{CANCELED}\}))$
$\mathtt{TL_{13}}$ A FAILED sub-task with fallbacks transitions to RETRY SCHEDULED, FALLBACK SELECTED, ERROR or is CANCELED.
$AG((state=\texttt{FAILED}\land has\_fallbacks)\rightarrow AX(state\in\{\texttt{RETRY},\texttt{FALLBACK},\texttt{ERROR},\texttt{CANCELED}\}))$
@@
5 Property Specification
The formalization of the HA and Task Lifecycle models provides the foundation for analyzing and verifying critical system properties. In this context, formal properties are verifiable statements about the system’s execution paths, used to ensure the safety, security, and functionality of an Agentic AI system’s behavior. We structure these properties into two distinct sets: those captured by the HA model (overall orchestration, Table 1), and those governing the task lifecycle model (sub-task state transitions, Table 2). These properties do not operate in isolation; they form a tightly coupled system of interdependent guarantees necessary for correct and reliable function in a multi-AI agent environment. Both sets of properties are expressed using temporal logic, CTL and LTL, which allow for an unambiguous definition of desired system behavior and execution paths.
5.1 Property Design Principles and Definitions
Both models include a set of property categories, each containing multiple specific properties, to establish verifiable guarantees. The core categories used in both tables are liveness, safety, completeness, fairness, and reachability. These properties ensure the overall system makes progress, avoids undesirable states, achieves its goals, and handles reliable resource allocation.
Liveness. Liveness ensures that an Agentic AI system will eventually make progress toward a goal. This property guarantees that, despite asynchronous actions, inter-agent dependencies, or potential conflicts, the system avoids global deadlock or infinite starvation. This is critical here, where a lack of progress in one agent can propagate and result in a cascading stall of the entire system.
  • Example ($\mathtt{HP_{1}}$): Every user prompt must eventually receive a final response from the host agent. This is expressed in temporal logic as $AG(Req_{U}\to AF\ Resp_{H})$.
  • Example ($\mathtt{TL_{1}}$): A sub-task that is CREATED must eventually terminate in a defined end state (COMPLETED, ERROR, or CANCELED).